# 2INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy 3Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 4Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany

### v1

@@ 4.2 Model selection with GW events
Here, we compare our models by calculating the detectability-conditioned posterior odds for GW events with BHNS components. To carry out this comparison, we follow the formalism presented in Mould et al. (2023), which we briefly summarize in Appendix C. Selection effects are computed using public LVK injections (Abbott et al. 2023) with a detection threshold in the false-alarm rate $<0.25\,\mathrm{yr}^{-1}$ (Abbott et al. 2023b).
Table 7 presents the detectability-conditioned Bayes factors, $\ln\mathcal{D}_{A/B}$, where model $A$ corresponds to a Pop. III model (as listed in Table 1) and model $B$ represents the fiducial Pop. II model (kro1). These values are obtained by convolving the merger catalogs generated with sevn with $\mathcal{R}$ate, following the methodology described in Sec. 2.7, and considering only mergers occurring at redshifts $z<2$. We compute Bayes factors for three GW events: GW190917, GW191219, and GW200105 (Abbott et al. 2021, 2024, 2023a). Our analysis includes GW190917, with a secondary mass within the lower mass gap ($2-5\,{\rm M}_{\odot}$). The waveform parameters used to evaluate the Bayes factors are the primary mass, mass ratio, and redshift of each event. In this analysis, we only consider events with sufficient overlap between the observed parameters and the simulated distributions to ensure meaningful Bayes factor estimates. A decisive preference between models is typically indicated when $|\ln\mathcal{D}_{A/B}|\gtrsim 5$ (Jeffreys 1939).
Among the selected GW events, only one (GW191219) has primary mass $m_{1}>20\,{\rm M}_{\odot}$. It is generally expected that Pop. II BHNS mergers with such high primary masses form in star clusters (Rastello et al. 2020; Arca Sedda 2020; Arca sedda et al. 2024). However, as illustrated in Figs. 5 and 6, an alternative formation channel could derive from the evolution of isolated Pop. III binaries.
GW191219 favors a Pop. III origin. This system is characterized by a highly asymmetric mass ratio ($q\sim 0.04$), leading to higher Bayes factors for models that favor mergers with extremely low mass ratios, such as kro5, lar5, log3, log4, log5, and top5. This outcome is consistent with Fig. 11, which shows that all simulated Pop. III models predominantly yield BHNS mergers with $q<0.1$. It should be noted that the false alarm rate associated with GW191219 is $\sim 4\,\rm{yr}^{-1}$, which is not consistent with the threshold typically adopted for estimating selection effects in the calculation of detection-conditioned Bayes factors. Nevertheless, the detectability-free Bayes factors $\mathcal{B}_{\rm Pop.\leavevmode\nobreak\ III/Pop.\leavevmode\nobreak\ II}$ for this event are found to be in close agreement with the detectability-weighted values reported in the table.
The remaining two GW events, GW190917 and GW200105, both have primary masses $m_{1}<20\,{\rm M}_{\odot}$ and show stronger support for a Pop. II formation channel.
As a consequence, we find that Pop. III BHNS mergers contribute to the population of systems with high primary mass and low mass ratio at low redshift more than Pop. II. In contrast, for $m_{1}<15\,{\rm M}_{\odot}$, Pop. II binaries remain the dominant formation channel.
@@
GW200105 $9.1_{-1.7}^{+1.7}$ $1.91_{-0.24}^{+0.33}$ $-4.4$ $-11.2$ $-4.7$ $-9.9$ $-2.7$ $-16$ $-3.7$ $-3$ $-14.8$ $-3.3$ $-512.4$
Table 7: Detectability-conditioned Bayes factors $\ln\mathcal{D}$ for Pop. III models, evaluated against the fiducial Pop. II model (kro1), for GW events classified as BHNS mergers or containing one component in the lower mass gap. The selected events were detected during the first three observing runs of LVK. Columns 1–3: event name and source-frame component masses. Columns 4–14: $\ln\mathcal{D}$ values for each Pop. III model relative to the Pop. II reference model.
@@ 5 Summary
We have investigated the properties of black hole–neutron star binary (BHNS) and binary neutron star (BNS) mergers from Pop. III and Pop. II stars, by means of binary population synthesis simulations performed with sevn (Costa et al. 2023). These merger catalogs encompass the uncertainties on initial mass functions (IMF) and orbital parameters of Pop. III stars.
We have estimated the merger rate densities (MRDs) of both BHNSs (Fig.3) and BNSs (Fig.8) from Pop. III stars and compared them to the MRDs of Pop. II stars.
@@
We find that the MRD of BHNSs from Pop. III stars ranges from $\sim 10^{-2}$ to $2\,\rm Gpc^{-3}\,yr^{-1}$ at redshift $z\sim{13}$. Models with initial orbital parameters from Stacy & Bromm (2013) produce smaller MRDs with respect to models with distributions from Sana et al. (2012). This behavior was already pointed out in Santoliquido et al. (2023) in the case of binary black hole (BH) mergers. The MRD of BHNSs from Pop. II stars dominates over the one from Pop. III stars at least up to $z\sim 17$.
The delay times $t_{\rm del}$ of BHNS mergers for our fiducial models (Fig. 4) follow quite closely the $\propto t^{-1}$ trend from Dominik et al. (2012).
@@
We compare the primary BH mass $m_{1}$ distribution of BHNS mergers for the fiducial models, and find that BHs from Pop. III stars tend to produce a flat distribution in $m_{1}$ with a peak around $22\,{\rm M}_{\odot}$, while BHs from Pop. II stars present a peak at $m_{1}<15\,{\rm M}_{\odot}$ and a decreasing trend up to $\sim 60\,{\rm M}_{\odot}$ (Fig. 5). These different trends in $m_{1}$ result in different trends in the mass ratio $q=m_{2}/m_{1}$ (Fig. 7). BHNSs from Pop. III stars yield a distribution peaking at $q<0.1$ and extending up to $q\sim 0.3$, whereas Pop. II BHNS mergers peak at $q\sim 0.2$ and extend up to $q\sim 0.35$.
@@
Pop. III high-mass BHNSs merge preferentially with large delay times $t_{\rm del}>10^{3}\,\rm Myr$ (Fig. 6). Consequently, second-generation interferometers might be able to detect BHNS mergers with $m_{1}>20\,{\rm M}_{\odot}$ from a metal-free population in the local Universe.
@@
BNSs from Pop. III and II stars mainly merge with short delay times ($t_{\rm del}<50\,\rm Myr$) because of their formation channels, which usually involve one or more common envelope events (Fig. 9). As a result, the peaks of the MRDs of BNSs from both Pop. III and Pop. II stars are close to the peaks of their respective star formation rate density models (Hartwig et al. 2022, 2024; Liu et al., in prep.). We find that the MRDs of metal-free BNSs range from $\sim 10^{-3}$ to $\sim 15\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ at $z\sim 15$. The MRD of Pop. II BNSs peaks at $\mathcal{R}_{\rm BNS}(z)\sim 5\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ around $z\sim 10$, dominating over the Pop. III contribution up to that redshift. Despite the compactness of Pop. III stars during their evolution (Fig. 1), Pop. II stars produce a MRD at least one order of magnitude higher than the Pop. III one. This is mainly due to the impact of the star formation rate density models on the MRDs. Also in this case, configurations with initial orbital parameters from Stacy & Bromm (2013) produce smaller MRDs with respect to those with initial distributions from Sana et al. (2012).
@@
We explore the impact of the initial conditions on the distributions of primary mass $m_{1}$ (Fig. 10) and mass ratio $q$ (Fig. 11) of BHNS mergers from Pop. III stars. We find that all simulated models produce merging BHNSs with $6\lesssim m_{1}\lesssim 50\,{\rm M}_{\odot}$, except for one (log3), which yields BHNS mergers with primary mass up to $400\,{\rm M}_{\odot}$. Models with initial orbital parameters from Stacy & Bromm (2013) create preferentially merging systems with $m_{1}>15\,{\rm M}_{\odot}$, while models with initial orbital parameters from Sana et al. (2012) yield flatter $m_{1}$ distributions with a peak around $22\,{\rm M}_{\odot}$. The impact of the initial orbital configuration is milder for Pop. II stars, because of their largest expansion during their evolution.
@@
We compute the detectability-conditioned Bayes factors (Mould et al. 2023) for three GW signals detected by LVK, classified either as BHNS mergers or as compact mergers with a secondary mass in the lower mass gap ($2-5\,{\rm M}_{\odot}$, Fryer et al. 2012). In our analysis, we compare each simulated Pop. III model to the fiducial Pop. II model (see Table 7). GW191219 stands out as favoring a Pop. III origin, particularly for models that predict mergers with extreme mass ratios. In contrast, GW signals with $m_{1}<20\,{\rm M}_{\odot}$ favor a Pop. II origin.
@@
In summary, our results indicate that, although Pop. III stars are more compact than Pop. II stars, their lower star formation rate density leads to MRDs of BHNSs and BNSs that are at least one order of magnitude lower than for BHNSs and BNSs born from Pop. II binaries. This trend depends on the chosen models for the star formation rate density of Pop. III (Hartwig et al. 2022) and Pop. II stars (Hartwig et al. 2024, Liu et al., in prep.). Finally, we find that Pop. III BHNS mergers can involve massive BHs, and provide a possible formation channel for BHNS mergers with BH mass $>20\,{\rm M}_{\odot}$ detected in the local Universe. The implications of our results for detection rates with second- and third-generation interferometers, as well as the effects of rotation and of different supernova models on the component masses of BHNS mergers, will be explored in future works.
@@ Acknowledgements.
MMa, GC, GI, and FS acknowledge financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017. MMa, RSK and BL also acknowledge financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. MB and MMa also acknowledge support from the PRIN grant METE under the contract no. 2020KB33TP. FS acknowledges financial support from the AHEAD2020 project (grant agreement n. 871158). MAS acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101025436 (project GRACE-BH) and from the MERAC Foundation. GC acknowledges partial financial support from European Union—Next Generation EU, Mission 4, Component 2, CUP: C93C24004920006, project ‘FIRES’. GI is supported by a fellowship grant from the la Caixa Foundation (ID 100010434), code LCF/BQ/PI24/12040020. GI also acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.4, - Call for tender No. 3138 of 18/12/2021 of Italian Ministry of University and Research funded by the European Union – NextGenerationEU. MMo is supported by the LIGO Laboratory through the National Science Foundation awards PHY-1764464 and PHY-2309200. RSK acknowledges financial support from the ERC via Synergy Grant ”ECOGAL” (project ID 855130) and from the German Ministry for Economic Affairs and Climate Action in project ”MAINN” (funding ID 50OO2206). RSK also thanks the 2024/25 Class of Radcliffe Fellows for highly interesting and stimulating discussions. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1597-1 FUGG and INST 35/1503-1 FUGG. We used sevn (https://gitlab.com/sevncodes/sevn) to generate our catalogs (Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023). We used $\mathcal{R}$ate (https://gitlab.com/Filippo.santoliquido/cosmo_rate_public) to generate our merger rate densities and convolve our catalogs of mergers with the star formation history (Santoliquido et al. 2020, 2021, 2023). We obtained our star formation rate densities from a-sloth (https://gitlab.com/thartwig/asloth). This research made use of NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), Pandas (pandas development team 2020). For the plots we used Matplotlib (Hunter 2007).
@@
In the following, we briefly present the initial distributions of primary mass, mass ratio, orbital period and eccentricity originally described by Costa et al. (2023) and listed in Table 1.
@@ A.1 Initial mass functions
    A Kroupa (2001) IMF $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-2.3}$, which is commonly adopted for Pop. II stars. With respect to the canonical IMF, here we consider a single slope since $m_{\textrm{min}}\geq 0.5\,{\rm M}_{\odot}$.
@@
    A Larson (1998) IMF $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-2.35}\exp{\left(-m_{\mathrm{cut1}}/m_{\mathrm{ZAMS}}\right)}$, with $m_{\mathrm{cut1}}=20\,{\rm M}_{\odot}$.
@@
    A flat log distribution $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-1}$ (Stacy & Bromm 2013; Hirano et al. 2015; Susa et al. 2014; Wollenberg et al. 2020; Chon et al. 2021; Tanikawa et al. 2021b; Jaura et al. 2022; Prole et al. 2022).
@@
    A top heavy distribution (Stacy & Bromm 2013; Jaacks et al. 2019; Liu & Bromm 2020a), $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-0.17}\exp{\left(-m_{\mathrm{cut2}}^{2}/m_{\mathrm{ZAMS}}^{2}\right)}$, where $m_{\mathrm{cut2}}=20\,M_{\odot}$.

### v2

@@ 4.2 Model selection with GW events
Here, we compare our models by calculating the detectability-conditioned posterior odds for GW events with BHNS components. To carry out this comparison, we followed the formalism presented in mould2023, which we briefly summarize in Appendix D. Selection effects were computed using public LVK injections (KAGRA:2023pio) with a detection threshold in the false-alarm rate of $0.25\,\mathrm{yr}^{-1}$ (abbottO3popandrate).
[…]
Among the selected GW events, only one (GW191219) has a primary mass of $m_{1}>20\,{\rm M}_{\odot}$. It is generally expected that Pop. II BHNS mergers with such high primary masses form in star clusters (rastello2020; mas2020; mas2024). However, as is illustrated in Figs. 5 and 6, an alternative formation channel could derive from the evolution of isolated Pop. III binaries.
GW191219 favors a Pop. III star origin. This system is characterized by a highly asymmetric mass ratio ($q\sim 0.04$), leading to higher Bayes factors for models that favor mergers with extremely low mass ratios, such as kro5, lar5, log3, log4, log5, and top5. This outcome is consistent with Fig. 11, which shows that all simulated Pop. III models predominantly yield BHNS mergers with $q<0.1$. It should be noted that the false alarm rate associated with GW191219 is $\sim 4\,\rm{yr}^{-1}$, which is not consistent with the threshold typically adopted for estimating selection effects in the calculation of detection-conditioned Bayes factors. Nevertheless, the detectability-free Bayes factors, $\mathcal{B}_{\rm Pop.\penalty 10000\ III/Pop.\penalty 10000\ II}$, for this event are found to be in close agreement with the detectability-weighted values reported in the table. The remaining two GW events, GW190917 and GW200105, both have primary masses of $m_{1}<20\,{\rm M}_{\odot}$ and show stronger support for a Pop. II formation channel compared to the Pop. III scenario.
As a consequence, we find that Pop. III BHNS mergers contribute to the population of systems with a high primary mass and a low mass ratio at low redshift more than Pop. II stars. In contrast, for $m_{1}<15\,{\rm M}_{\odot}$, Pop. II binaries remain the dominant formation channel for the low-redshift BHNS mergers.
Table 7: Detectability-conditioned Bayes factors, $\ln\mathcal{D}$, for Pop. III models against the fiducial Pop. II model [^7] Detectability-conditioned Bayes factors, $\ln\mathcal{D}$, for Pop. III models, evaluated against the fiducial Pop. II model (kro1), for GW events classified as BHNS mergers or containing one component in the lower mass gap. The selected events were detected during the first three observing runs of LVK. Columns 1–3: Event name and source-frame component masses. Columns 4–14: $\ln\mathcal{D}$ values for each Pop. III model relative to the Pop. II reference model.
@@
GW200105 $9.1_{-1.7}^{+1.7}$ $1.91_{-0.24}^{+0.33}$ 4.4-4.4 11.2-11.2 4.7-4.7 9.9-9.9 2.7-2.7 16-16 3.7-3.7 3-3 14.8-14.8 3.3-3.3 512.4-512.4
4.3 Caveats and future prospects
As is shown in Table 1, we have explored a wide range of Pop. III IMFs and orbital parameter models. However, our results are still affected by several uncertainties in stellar and binary evolution.
Stellar structure assumptions are a key source of uncertainty. For instance, suppressing core overshooting leads to more compact stars, which enhances the merger rate and increases the number of systems forming BHs above the mass gap (tanikawa2021; tanikawa2021b; tanikawa2022; tanikawa2023; tanikawa2024). Moreover, Pop. III stars are expected to be rapid rotators, potentially undergoing chemically homogeneous evolution. As is shown by santoliquido2023 and mestichelli2024, this channel produces very compact stars, again boosting merger rates and potentially leading to BHNS mergers with primary masses above the upper mass gap.
Another uncertainty concerns the initial distributions of the orbital period, mass ratio, and eccentricity. Current hydrodynamical simulations are limited and typically biased toward longer orbital periods (e.g., sb2013; sugimura2020; park2023; klessen2023). To account for this, we adopt two prescriptions for orbital parameter distributions, from sana2012 and sb2013, chosen to span a plausible range of outcomes. We also note that the distributions of sana2012, derived for local O- and B-type stars, have already been employed in several recent studies (tanikawa2022; tanikawa2023; costa2023; santoliquido2023; tanikawa2024; mestichelli2024).
Supernova physics represents another major source of uncertainty. Throughout this work we have adopted the rapid SN model of fryer2012. Alternative prescriptions could alter the component masses and mass-ratio distributions of compact-object binaries. For instance, the delayed SN model of fryer2012 allows neutron-star masses to extend up to $3\,{\rm M}_{\odot}$ and BH masses down to the same value. This would shift the resulting mass-ratio distribution toward higher values compared to Fig. 7. We shall explore the impact of SN models on BNS and BHNS mergers from Pop. III in future work. In addition, recent studies (disberg2025a; disberg2025b) indicate that the natal kicks derived by hobbs2005 are likely overestimated by $\sim 50\%$, due to an erroneous histogram representation. Correcting for this yields systematically lower kicks, which increases the survival probability of binaries containing neutron stars. Nevertheless, even the revised kick distributions remain sufficiently strong to disrupt a large fraction of systems, particularly those with wide orbits. Updated models for SN natal kicks will be taken into account as well for future work.
Finally, in this work we have focused exclusively on Pop. III and early Pop. II stars. This choice was motivated by our goal of investigating their role at high redshift, where they are expected to be most relevant, and in view of the future observations that will be conducted by third-generation GW interferometers. We have chosen to also keep this work focused on metal-poor populations in the case of the model selection performed in Sec. 4.2.
@@ 5 Summary
We have investigated the properties of BHNS and BNS mergers from Pop. III and Pop. II stars, by means of binary population synthesis simulations performed with sevn (costa2023). These merger catalogs encompass the uncertainties on IMFs and orbital parameters of Pop. III stars. We have estimated the merger rate densities (MRDs) of both BHNSs (Fig. 3) and BNSs (Fig. 8) from Pop. III stars and compared them to the MRDs of Pop. II stars.
@@
We find that the MRD of BHNSs from Pop. III stars ranges from $\sim 10^{-2}$ to $2\,\rm Gpc^{-3}\,yr^{-1}$ at redshift $z\sim{13}$. Models with initial orbital parameters from sb2013 produce smaller MRDs with respect to models with distributions from sana2012. This behavior was already pointed out in santoliquido2023 in the case of binary BH mergers. The MRD of BHNSs from Pop. II stars dominates over the one from Pop. III stars at least up to $z\sim 17$. The delay times, $t_{\rm del}$, of BHNS mergers for our fiducial models (Fig. 4) follow the $\propto t^{-1}$ trend from dominik2012 quite closely.
@@
We compare the primary BH mass, $m_{1}$, distribution of BHNS mergers for the fiducial models, and find that BHs from Pop. III stars tend to produce a flat distribution in $m_{1}$ with a peak around $22\,{\rm M}_{\odot}$, while BHs from Pop. II stars present a peak at $m_{1}<15\,{\rm M}_{\odot}$ and a decreasing trend up to $\sim 60\,{\rm M}_{\odot}$ (Fig. 5). These different trends in $m_{1}$ result in different trends in the mass ratio, $q=m_{2}/m_{1}$ (Fig. 7). The BHNSs from Pop. III stars yield a distribution peaking at $q<0.1$ and extending up to $q\sim 0.3$, whereas Pop. II BHNS mergers peak at $q\sim 0.2$ and extend up to $q\sim 0.35$.
@@
Population III high-mass BHNSs merge preferentially with large delay times, $t_{\rm del}>10^{3}\,\rm Myr$ (Fig. 6). Consequently, second-generation interferometers might be able to detect BHNS mergers with $m_{1}>20\,{\rm M}_{\odot}$ from a metal-free population in the local Universe.
@@
The BNSs from Pop. III and II stars mainly merge with short delay times ($t_{\rm del}<50\,\rm Myr$) because of their formation channels, which usually involve one or more common envelope events (Fig. 9). As a result, the peaks of the MRDs of BNSs from both Pop. III and Pop. II stars are close to the peaks of their respective star formation rate density models (hartwig2022; hartwig2024; liu2025). We find that the MRDs of metal-free BNSs range from $\sim 10^{-3}$ to $\sim 15\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ at $z\sim 15$. The MRD of Pop. II BNSs peaks at $\mathcal{R}_{\rm BNS}(z)\sim 5\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ around $z\sim 10$, dominating over the Pop. III contribution up to that redshift. Despite the compactness of Pop. III stars during their evolution (Fig. 1), Pop. II stars produce a MRD at least one order of magnitude higher than the Pop. III one. This is mainly due to the impact of the star formation rate density models on the MRDs. Also in this case, configurations with initial orbital parameters from sb2013 produce smaller MRDs with respect to those with initial distributions from sana2012.
@@
We explored the impact of the initial conditions on the distributions of the primary mass, $m_{1}$ (Fig. 10), and mass ratio, $q$ (Fig. 11), of BHNS mergers from Pop. III stars. We find that all simulated models produce merging BHNSs with $6\lesssim m_{1}\lesssim 50\,{\rm M}_{\odot}$, except for one (log3), which yields BHNS mergers with primary masses of up to $400\,{\rm M}_{\odot}$. Models with initial orbital parameters from sb2013 create preferentially merging systems with $m_{1}>15\,{\rm M}_{\odot}$, while models with initial orbital parameters from sana2012 yield flatter $m_{1}$ distributions with a peak around $22\,{\rm M}_{\odot}$. The impact of the initial orbital configuration is milder for Pop. II stars, because of their largest expansion during their evolution.
@@
We computed the detectability-conditioned Bayes factors (mould2023) for three GW signals detected by LVK, classified either as BHNS mergers or as compact mergers with a secondary mass in the lower-mass gap ($2-5\,{\rm M}_{\odot}$, fryer2012). In our analysis, we compared each simulated Pop. III model to the fiducial Pop. II model (see Table 7). GW191219 stands out as favoring a Pop. III origin, particularly for models that predict mergers with extreme mass ratios. In contrast, GW signals with $m_{1}<20\,{\rm M}_{\odot}$ favor a Pop. II origin.
@@
In summary, our results indicate that, although Pop. III stars are more compact than Pop. II stars, their lower star formation rate density leads to MRDs of BHNSs and BNSs that are at least one order of magnitude lower than for BHNSs and BNSs born from Pop. II binaries. This trend depends on the chosen models for the star formation rate density of Pop. III (hartwig2022) and Pop. II stars (hartwig2024; liu2025). Finally, we find that Pop. III BHNS mergers can involve massive BHs, and provide a possible formation channel for BHNS mergers with BH mass $20\,{\rm M}_{\odot}$ detected in the local Universe. The implications of our results for detection rates with second- and third-generation interferometers, as well as the effects of rotation and of different SN models and natal kick prescriptions on the component masses of BHNS mergers, will be explored in future works.
@@ Acknowledgements.
[…]
@@
In the following, we briefly present the initial distributions of primary mass, mass ratio, orbital period and eccentricity originally described by costa2023 and listed in Table 1.
@@ A.1 Initial mass functions
    A kroupa2001 IMF $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-2.3}$, which is commonly adopted for Pop. II stars. With respect to the canonical IMF, here we consider a single slope since $m_{\textrm{min}}\geq 0.5\,{\rm M}_{\odot}$.
@@
    A larson1998 IMF $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-2.35}\exp{\left(-m_{\mathrm{cut1}}/m_{\mathrm{ZAMS}}\right)}$, with $m_{\mathrm{cut1}}=20\,{\rm M}_{\odot}$.
@@
    A flat log distribution $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-1}$ (sb2013; hirano2015; susa2014; wollenberg2020; chon2021; tanikawa2021; jaura2022; prole2022).
@@
    A top heavy distribution (sb2013; jaacks2019; liu2020), $\xi(m_{\mathrm{ZAMS}})\propto m_{\mathrm{ZAMS}}^{-0.17}\exp{\left(-m_{\mathrm{cut2}}^{2}/m_{\mathrm{ZAMS}}^{2}\right)}$, where $m_{\mathrm{cut2}}=20\,M_{\odot}$.